OpenStudy (anonymous):

factor the expression 3 2x + 54 = 0 (the x is cubed)

OpenStudy (amistre64):

Factor 2x^3 +54 =0. x^3 = -54/2 x^3 = -27 cbrt(x^3) = cbrt(-27) x*x*x = (-3)*(-3)*(-3) x= -3 Double check: 2(-27) +54 = 0 -54 + 54 = 0 0=0 (its all good)

OpenStudy (anonymous):

ok yeah that makes sense thank you !!

OpenStudy (anonymous):

so then 2x 3 - 250 = 0 is same x 3 = 250/2 which is 125 x 3 = 125 so x = 5

OpenStudy (amistre64):

Yes, exactly :)

OpenStudy (anonymous):

ok thank you so much then